OpenStudy (anonymous):

Trevor is tiling his bathroom floor, which has an area that is represented as 120r3 square inches. Each tile has an area of square root of the quantity 16 r to the ninth power. The total number of tiles used can be represented by the expression below. one hundred twenty r to the third power, all over the square root of the quantity sixteen r to the ninth power.

OpenStudy (anonymous):

Let N denote the total number of tiles used N = (total area)/(area of one tile) = 175r^5/√(64r^17)

OpenStudy (anonymous):

could you explain that a little more please and thank you so much

OpenStudy (anonymous):

Simplify the denominator: √(64r^17) = √(64) √(r^17) = 8 √(r^16 * r^1) = 8 √(r^16) √r = 8 r^8 √r

OpenStudy (anonymous):

Putting this into the equation for N, N = 175r^5 / (8 r^8 √r) = (175/8) / (r^3 √r) = (175/8) / (r^3 r^½) = (175/8) /r^(7/2)

OpenStudy (anonymous):

which may also be written = (175/8) r^(-7/2)

OpenStudy (anonymous):

thank you!!!!!!!!

OpenStudy (anonymous):

medal please ?